OpenStudy (anonymous):

Solve. -20 = 5/6 x HAAAAAAALPPPPPPPP XD

OpenStudy (error1603):

HAAAAAAAAPPPPPPP! You watch him?...... No?.....Okay.

OpenStudy (anonymous):

lol

OpenStudy (error1603):

I'll help

OpenStudy (anonymous):

yay :)

OpenStudy (error1603):

Sorry its taking a long time im working on it

OpenStudy (texaschic101):

divide both sides by 5/6...or multiply both sides by 6/5

OpenStudy (anonymous):

It's fine ^-^

OpenStudy (error1603):

Gosh

OpenStudy (anonymous):

o3o i got - 100/60 and 25/36

OpenStudy (error1603):

So close to having it done.

OpenStudy (anonymous):

i got -100/120 ans 255/36 now o3o wat

OpenStudy (anonymous):

this is really confusing lol

OpenStudy (texaschic101):

- 20 * 6/5 = - 120/5 = - 24......x = 24

OpenStudy (anonymous):

oh thank you!

OpenStudy (anonymous):

^-^